The heights of WNBA basketball players are normally distributed. What percentage of these heights is within 2 standard deviations of the mean? 



1
Expert's answer
2021-06-07T18:39:31-0400

Use the 68–95–99.7 rule

Let "X" be a height "X\sim N(\mu, \sigma^2)," where "\ \mu" is the mean of the distribution, and "\sigma" is its standard deviation:


"P(\mu-2\sigma\leq X\leq \mu+2\sigma)\approx0.9545"

"P(\mu-2\sigma\leq X\leq \mu+2\sigma)"

"=P( X\leq \mu-2\sigma)-P( X< \mu+2\sigma)"

"=P( Z\leq \dfrac{\mu+2\sigma-\mu}{\sigma})-P( Z< \dfrac{\mu-2\sigma-\mu}{\sigma})"

"=P( Z\leq 2)-P( Z< -2)"

"\approx0.977249868-0.022750132\approx0.9545"